# AnyECG: Foundational Models for Electrocardiogram Analysis

### v1

@@ 3.1 ECG Datasets
To evaluate the performance of AnyECG and baseline models, we utilized a comprehensive set of ECG datasets that include all available unlabeled data during pretraining. These datasets cover a wide spectrum of cardiac conditions, patient demographics, and recording scenarios, ensuring robust testing across diverse settings. For various downstream tasks, we mixed all datasets together to minimize biases introduced by individual datasets and to better validate the model’s generalizability. This approach reduces the discrepancies arising from different data sources and enhances the unified capability of the model. The detailed data construction of the datasets can be found in Table 1. All datasets are formatted in WFDB format, including associated binary and text files that detail signal attributes and clinical annotations using SNOMED-CT codes. Detailed information about the datasets is provided in Appendix 7.1.
Table 1: Summary of ECG Datasets
@@
Dataset Recordings Sampling Rate Duration Notes CPSC [13] 6877 500 Hz 6–60 s Balanced sex CPSC-Extra [13] 3453 500 Hz 6–60 s Balanced sex INCART [14] 74 257 Hz 30 min each High-res; arrhythmia PTB [15] 516 1000 Hz Varies Wide range of pathologies PTB-XL [16] 21837 500 Hz 10 s Extensive clinical ECGs G12EC 10344 500 Hz Varies The Southeast’s unique demographics Undisclosed Dataset 10000 500 Hz 6–60 s Geographically distinct test set
3.2 Experimental Setup
Model Configurations.
@@ 3.2 Experimental Setup
Training Environment.
The pre-training of AnyECG was conducted on a comprehensive dataset compiled from seven different sources. For the downstream tasks, data splitting followed standard procedures, dividing the data into training and validation subsets using an 80/20 ratio. Binary cross-entropy loss was employed for binary classification tasks, while cross-entropy loss was utilized for multi-class classification tasks. Evaluation metrics for the downstream tasks are detailed in the Appendix 7.2. All experiments were executed on a computing cluster equipped with eight high-performance GPUs. We used the Adam optimizer with a learning rate of 1e-4 for all models training. Model selection was based on the best performance on validation sets, and final evaluations were conducted on separate test sets. To ensure the reliability of our results, performance metrics—including averages and standard deviations—were reported across five random seeds.
3.3 Experimental Results
Anomaly Detection.
In anomaly detection, we define normal category data as positive samples, with all other categories treated as anomalous samples. Table 2 compares AnyECG to state-of-the-art models in the anomaly detection task. AnyECG consistently outperforms other advanced models across all evaluation metrics. Specifically, the largest variant, AnyECG-XL, achieves the highest scores in accuracy, AUC-PR, AUROC, and Weighted F1 Score, demonstrating its strong ability to capture ECG signal characteristics. In contrast, traditional models like DENS-ECG [17] and ContraWR [18] show lower performance. DENS-ECG achieves moderate scores in accuracy and Weighted F1 Score, while ContraWR falls short in both metrics. Even the smaller versions of AnyECG, such as AnyECG-B and AnyECG-L, perform competitively and surpass most baseline models. This indicates that AnyECG maintains high performance across different scales without requiring extensive model parameters. Notably, the finetuned ECG-FM model [19] performs at an intermediate to above-average level compared to the baseline. However, as a pre-trained model, its performance may still be hindered by substantial differences between the pre-training data and the downstream task dataset, which likely impedes its ability to fully converge.
Table 2: Results Comparison with State-of-the-Art Models in Anomaly Detection
@@ 3.3 Experimental Results
Arrhythmia Detection.
Table 3 presents a performance comparison between AnyECG and other leading models in arrhythmia detection. The results show that AnyECG, particularly the AnyECG-XL variant, consistently outperforms competing models across all metrics. This demonstrates its strong ability to handle arrhythmia detection effectively. In contrast, models like DENS-ECG [17] and ContraWR [18] exhibit lower performance. Notably, although ECG-FM [19] employs pretraining, it achieves significantly lower accuracy. This underscores AnyECG’s robustness, as its consistent performance across all metrics confirms its suitability for real-world arrhythmia detection.
Table 3: Results Comparison with State-of-the-Art Models in Arrhythmia Detection
@@
Corrupted Lead Generation.
We evaluated AnyECG against CGAN [26] and WGAN [27] in generating corrupted ECG leads (see Table 4 and Figure 2). Using metrics like PSNR, SSIM, and MAE, AnyECG-L achieved the highest PSNR (32.7372 dB) and SSIM (0.8738), outperforming both CGAN and WGAN. Smaller models like AnyECG-L and AnyECG-B offer a better balance between capacity and generalization compared to AnyECG-XL. Due to limitations in its model architecture, ECG-FM [19] could not be applied to this task. Although the AnyECG models did not achieve the lowest MAE, this may be because they prioritize capturing abstract rhythms and morphological patterns over minimizing pixel-level errors in detailed, noisy signals. This suggests that while AnyECG effectively captures the overall structure and rhythm of ECG signals, it is somewhat less precise in reproducing finer details. Figure 2 shows the ECG signals generated by WGAN, CGAN, and AnyECG. Both WGAN and CGAN can capture general morphology but fail to accurately reproduce certain rhythms, leading to unsuccessful signal generation in those cases. AnyECG leverages two stage pre-training to capture complex rhythmic features, resulting in morphology closer to the original signals. However, it lacks detailed feature extraction in finer wave bands, leading to poorer reconstruction in these regions and higher MAE. These observations suggest that while AnyECG excels in preserving overall rhythmic and morphological integrity, there is room for improvement in reconstructing fine-grained details.
Table 4: Results Comparison with State-of-the-Art Models in Corrupted Lead Generation.
@@
Ultra-Long ECG Recognition.
@@
Recognizing ultra-long ECG signals is challenging due to their extended duration, rhythm variability, and noise, which require models to be robust and generalizable. Traditional time series models often struggle with high computational complexity, memory constraints, and difficulty in capturing long-term dependencies. Therefore, We proposed a hierarchical modeling approach that adapts to ultra-long ECG data by employing a sliding window method. As shown in Table 5, AnyECG, particularly the AnyECG-XL, achieves the highest scores across all evaluation metrics. This demonstrates its superior ability to capture complex patterns and maintain high accuracy when analyzing ultra-long ECG signals. Compared to state-of-the-art models like Inception1D [24] and RNN1D [22], AnyECG-XL shows a clear advantage, especially in AUROC and AUC-PR. Even the smaller variants, AnyECG-B and AnyECG-L, outperform most baseline models, highlighting AnyECG’s adaptability and scalability. The absence of results for the other pretrained ECG foundation model ECG-FM [19] is due to its inability to handle ultra-long sequence data, making it unsuitable for this task. In contrast, AnyECG’s consistent performance across all scales confirms its effectiveness in capturing key features of ultra-long ECG signals.
Table 5: Results Comparison with State-of-the-Art Models in Ultra-Long ECG Recognition
4 Related Works
Heterogeneous ECG Signal Analysis and Classification.
Self-supervised ECG Representation Learning.
5 Discussion
Social Impacts.
@@ 5 Discussion
6 Conclusion
In this paper, we proposed AnyECG, a foundation model family that learns universal embeddings through a two-stage self-supervised pre-training on seven diverse ECG datasets. AnyECG effectively handles the heterogeneity of ECG data through the design of a novel ECG Tokenizer, which includes a rhythm codebook and a multi-view synergistic decoder to learn representations from different proxy tasks. Additionally, the masked modeling in the second-stage pre-training plays a crucial role in enabling effective representation learning of both temporal and lead features of ECG signals. We validated various sizes of AnyECG models on multiple downstream tasks, including anomaly detection, arrhythmia detection, ECG lead generation, and ultra-long ECG signal recognition. Our experiments demonstrate that AnyECG outperforms all state-of-the-art methods in their respective fields, highlighting its effectiveness and versatility in ECG signal analysis.

### v2

@@ IV-A ECG Datasets
To evaluate the performance of AnyECG and baseline models, we utilized a comprehensive set of ECG datasets that include all available unlabeled data during pre-training. These datasets cover a wide spectrum of cardiac conditions, patient demographics, and recording scenarios, ensuring robust testing across diverse settings. For various downstream tasks, we mixed all datasets together to minimize biases introduced by individual datasets and to better validate the model’s generalizability. This approach reduces the discrepancies arising from different data sources and enhances the unified capability of the model. The detailed data construction of the datasets can be found in Table I. All datasets are formatted in WFDB format, including associated binary and text files that detail signal attributes and clinical annotations using SNOMED-CT codes.
To evaluate the performance of AnyECG and baseline models, we utilized a comprehensive set of ECG datasets that cover a wide spectrum of cardiac conditions, patient demographics, and recording scenarios, ensuring robust testing across diverse settings. The datasets include:
CPSC and CPSC-Extra Databases [31]: These consist of 12-lead ECG recordings ranging from 6 to 60 seconds in duration, sampled at 500 Hz, and include a balanced mix of male and female subjects.
INCART Database [32]: This database provides 74 annotated recordings extracted from 32 Holter records, each 30 minutes long and sampled at 257 Hz, offering high-resolution data ideal for arrhythmia classification.
PTB [33] and PTB-XL Databases [34]: The PTB Diagnostic ECG Database includes 516 recordings sampled at 1000 Hz, while PTB-XL contains 21,837 12-lead ECGs sampled at 500 Hz, each 10 seconds long, encompassing a wide range of cardiac pathologies.
Georgia 12-lead ECG Challenge (G12EC) Database: Comprising 10,344 recordings from the Southeastern United States, sampled at 500 Hz, this dataset adds demographic diversity to our evaluation.
Undisclosed Database: This dataset contributes an additional 10,000 ECG recordings, providing a geographically distinct test set to further validate the model’s performance without data leakage.
By employing this diverse collection of datasets, we thoroughly assess AnyECG’s ability to generalize across different patient populations, signal qualities, and clinical conditions.
IV-B Evaluation Metrics
We conducted four distinct experiments, each utilizing a specific set of evaluation metrics tailored to the task:
1. Anomaly Detection: Evaluated using Accuracy, AUC-PR (Area Under the Precision-Recall Curve), AUROC (Area Under the Receiver Operating Characteristic Curve), and Weighted F1 Score. These metrics assess the model’s ability to correctly identify anomalies and handle class imbalances effectively.
2. Arrhythmia Detection: Assessed with Accuracy, AUC-PR, Weighted F1 Score, and Precision. This combination of metrics evaluates the model’s performance in detecting various types of arrhythmias, emphasizing both overall accuracy and the precision of positive predictions.
3. Corrupted Lead Generation: Measured using PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity Index), and MAE (Mean Absolute Error). These metrics quantify the quality of the generated ECG signals by comparing the reconstructed signals to the original ones, focusing on signal fidelity and structural similarity.
4. Ultra-Long ECG Analysis: Evaluated using Accuracy, AUC-PR, AUROC, and Weighted F1 Score. These metrics measure the model’s effectiveness in accurately identifying anomalies and handling class imbalances.
The definitions of the evaluation metrics used across these experiments are as follows: Accuracy: The proportion of correctly predicted instances out of all instances, indicating the overall effectiveness of the model. Precision: The ratio of true positive predictions to the total number of positive predictions, reflecting the model’s ability to avoid false positives. AUC-PR (Area Under the Precision-Recall Curve): Measures the trade-off between precision and recall for different threshold settings, particularly useful for imbalanced datasets. AUROC (Area Under the Receiver Operating Characteristic Curve): Represents the model’s ability to distinguish between classes across all classification thresholds. Weighted F1 Score: The harmonic mean of precision and recall, weighted by the number of true instances for each class, providing a balanced evaluation of the model’s performance. PSNR (Peak Signal-to-Noise Ratio): Indicates the quality of signal reconstruction by comparing the maximum possible signal power to the power of reconstruction noise, with higher values signifying better quality. SSIM (Structural Similarity Index): Assesses the similarity between two signals in terms of luminance, contrast, and structure, with values closer to 1 indicating higher similarity. MAE (Mean Absolute Error): Represents the average absolute difference between predicted and actual values, serving as a measure of prediction accuracy. By employing these tailored metrics across different experiments, we ensure a comprehensive evaluation of our model’s performance in various aspects of ECG signal processing and classification.
TABLE I: Summary of ECG Datasets
@@ IV-B Evaluation Metrics
Dataset Recordings Sampling Rate Duration Notes CPSC [31] 6877 500 Hz 6–60 s Balanced sex CPSC-Extra [31] 3453 500 Hz 6–60 s Balanced sex INCART [32] 74 257 Hz 30 mins High-res; arrhythmia PTB [33] 516 1000 Hz Varies Wide range of pathologies PTB-XL [34] 21837 500 Hz 10 s Extensive clinical ECGs G12EC 10344 500 Hz Varies The Southeast’s unique demographics Undisclosed Dataset 10000 500 Hz 6–60 s Geographically distinct test set
IV-C Experimental Setup
Model Configurations
@@ IV-C Experimental Setup
Training Environment
The pre-training of AnyECG was conducted on a comprehensive dataset compiled from seven different sources. For the downstream tasks, data splitting followed standard procedures, dividing the data into training and validation subsets using an 80/20 ratio. Binary cross-entropy loss was employed for binary classification tasks, while cross-entropy loss was utilized for multi-class classification tasks. All experiments were executed on a computing cluster equipped with eight high-performance GPUs. We used the Adam optimizer with a learning rate of 1e-4 for all models training. Model selection was based on the best performance on validation sets, and final evaluations were conducted on separate test sets. To ensure the reliability of our results, performance metrics—including averages and standard deviations—were reported across five random seeds.
IV-D Experimental Results
Anomaly Detection
In anomaly detection, we define normal category data as positive samples, with all other categories treated as anomalous samples. Table II compares AnyECG to state-of-the-art models in the anomaly detection task. AnyECG consistently outperforms other advanced models across all evaluation metrics. Specifically, the largest variant, AnyECG-XL, achieves the highest scores in accuracy, AUC-PR, AUROC, and Weighted F1 Score, demonstrating its strong ability to capture ECG signal characteristics. In contrast, traditional models like DENS-ECG [35] and ContraWR [36] show lower performance. DENS-ECG achieves moderate scores in accuracy and Weighted F1 Score, while ContraWR falls short in both metrics. Even the smaller versions of AnyECG, such as AnyECG-B and AnyECG-L, perform competitively and surpass most baseline models. This indicates that AnyECG maintains high performance across different scales without requiring extensive model parameters. Notably, the finetuned ECG-FM model [37] performs at an intermediate to above-average level compared to the baseline. However, as a pre-trained model, its performance may still be hindered by substantial differences between the pre-training data and the downstream task dataset, which likely impedes its ability to fully converge.
TABLE II: Results Comparison with State-of-the-Art Models in Anomaly Detection
@@ IV-D Experimental Results
TABLE III: Results Comparison with State-of-the-Art Models in Ultra-Long ECG Analysis
Arrhythmia Detection
Table IV presents a performance comparison between AnyECG and other leading models in arrhythmia detection. The results show that AnyECG, particularly the AnyECG-XL variant, consistently outperforms competing models across all metrics. This demonstrates its strong ability to handle arrhythmia detection effectively. In contrast, models like DENS-ECG [35] and ContraWR [36] exhibit lower performance. Notably, although ECG-FM [37] employs pre-training, it achieves significantly lower accuracy. This underscores AnyECG’s robustness, as its consistent performance across all metrics confirms its suitability for real-world arrhythmia detection.
TABLE IV: Results Comparison with State-of-the-Art Models in Arrhythmia Detection
@@
Corrupted Lead Generation
We evaluated AnyECG against CGAN [44] and WGAN [45] in generating corrupted ECG leads (see Table V and Figure 3). Using metrics like PSNR, SSIM, and MAE, AnyECG-L achieved the highest PSNR (32.7372 dB) and SSIM (0.8738), outperforming both CGAN and WGAN. Smaller models like AnyECG-L and AnyECG-B offer a better balance between capacity and generalization compared to AnyECG-XL. Due to limitations in its model architecture, ECG-FM [37] could not be applied to this task. Although the AnyECG models did not achieve the lowest MAE, this may be because they prioritize capturing abstract rhythms and morphological patterns over minimizing pixel-level errors in detailed, noisy signals. This suggests that while AnyECG effectively captures the overall structure and rhythm of ECG signals, it is somewhat less precise in reproducing finer details. Figure 3 shows the ECG signals generated by WGAN, CGAN, and AnyECG. Both WGAN and CGAN can capture general morphology but fail to accurately reproduce certain rhythms, leading to unsuccessful signal generation in those cases. AnyECG leverages two stage pre-training to capture complex rhythmic features, resulting in morphology closer to the original signals. However, it lacks detailed feature extraction in finer wave bands, leading to poorer reconstruction in these regions and higher MAE. These observations suggest that while AnyECG excels in preserving overall rhythmic and morphological integrity, there is room for improvement in reconstructing fine-grained details.
TABLE V: Results Comparison with State-of-the-Art Models in Corrupted Lead Generation.
@@
Ultra-Long ECG Analysis
@@
Recognizing ultra-long ECG signals is challenging due to their extended duration, rhythm variability, and noise, which require models to be robust and generalizable. Traditional time series models often struggle with high computational complexity, memory constraints, and difficulty in capturing long-term dependencies. Therefore, We proposed a hierarchical modeling approach that adapts to ultra-long ECG data by employing a sliding window method. As shown in Table III, AnyECG, particularly the AnyECG-XL, achieves the highest scores across all evaluation metrics. This demonstrates its superior ability to capture complex patterns and maintain high accuracy when analyzing ultra-long ECG signals. Compared to state-of-the-art models like Inception1D [42] and RNN1D [40], AnyECG-XL shows a clear advantage, especially in AUROC and AUC-PR. Even the smaller variants, AnyECG-B and AnyECG-L, outperform most baseline models, highlighting AnyECG’s adaptability and scalability. The absence of results for the other pretrained ECG foundation model ECG-FM [37] is due to its inability to handle ultra-long sequence data, making it unsuitable for this task. In contrast, AnyECG’s consistent performance across all scales confirms its effectiveness in capturing key features of ultra-long ECG signals.
IV-E Pre-training Phase Ablation Study
To evaluate the contribution of each component in our pre-training strategy, we conducted an ablation study focusing on the pre-training phases. Specifically, we analyzed the effects of pre-training the ECG Tokenizer and the full AnyECG foundation model on anomaly detection performance. Table VI and Figure 4 presents the results of this study. The first configuration is AnyECG-B without ECG Tokenizer pre-training. The second configuration includes a pre-trained ECG Tokenizer but skips pre-training the AnyECG foundation model. The final configuration involves full pre-training of both the ECG Tokenizer and the AnyECG. The results show that pre-training the ECG Tokenizer leads to noticeable improvements over the baseline. This indicates that a pre-trained ECG Tokenizer enhances the model’s ability to capture meaningful representations of the ECG signals. When the full AnyECG foundation model is also pre-trained, we observe a significant performance boost across all metrics. These gains underscore the importance of comprehensive pre-training in enhancing the model’s anomaly detection capabilities. The fact that full pre-training yields the best results confirms that both components—the ECG Tokenizer and the AnyECG—contribute positively to the overall performance. Pre-training the AnyECG foundation model allows it to learn generalizable features that are beneficial for downstream tasks, while the pre-trained ECG Tokenizer ensures effective encoding of the input signals.
Figure 4: Ablation study of Pre-training Phase in Anomaly Detection with AnyECG-L and AnyECG-XL
TABLE VI: Ablation study of Pre-training Phase in Anomaly Detection
Methods Accuracy $\uparrow$ AUC-PR $\uparrow$ AUROC $\uparrow$ Weighted F1 Score $\uparrow$ AnyECG-B 0.7623 0.9243 0.7729 0.7512 w/o ECG Tokenizer AnyECG-B 0.7810 0.9358 0.8232 0.7826 w/o Pre-train AnyECG AnyECG-B 0.8188 0.9517 0.8502 0.8863
TABLE VII: Ablation Study on Loss Function Components in Anomaly Detection
Loss Configuration Accuracy $\uparrow$ AUC-PR $\uparrow$ AUROC $\uparrow$ Weighted F1 Score $\uparrow$ Full Loss 0.8188 0.9517 0.8502 0.8863 w/o Morphology Loss 0.8059 0.9373 0.8381 0.8754 w/o Frequency Loss 0.8125 0.9412 0.8475 0.8621 w/o Demography Loss 0.8134 0.9487 0.8445 0.8801 w/o Codebook Loss 0.7522 0.8950 0.7900 0.8150 w/o Commitment Loss 0.7855 0.9275 0.8225 0.8575
IV-F Loss Function Ablation Study
To assess the effectiveness of our loss function design, we conducted an ablation study by systematically removing each component of the loss function. This allowed us to evaluate how each term contributes to the model’s ability to capture meaningful features from ECG signals. Table VII presents the results of this study. The ”Full Loss” configuration, which includes all components of our proposed loss function, serves as the baseline for optimal performance. When we individually removed each loss component, we observed a decrease in performance across all evaluation metrics. Omitting the Morphology Loss resulted in a noticeable decline, indicating its significant role in helping the model capture the morphological characteristics of ECG signals, which are crucial for accurate anomaly detection. Excluding the Frequency Loss also led to reduced performance, suggesting that capturing frequency domain information is important for understanding underlying patterns in ECG signals. Removing the Demography Loss caused a performance drop as well, though to a lesser extent compared to the Morphology and Frequency losses. This highlights that incorporating demographic information refines the model’s predictions by accounting for variations in ECG patterns across different demographic groups. The most significant decrease in performance was observed when the Codebook Loss was removed. This component is essential for encouraging diversity and utilization of codebook entries in the vector quantization process, playing a critical role in the model’s ability to represent ECG signals effectively. Lastly, removing the Commitment Loss also led to a decline in performance, though the impact was less severe than omitting the Codebook Loss. The Commitment Loss ensures consistency in the representation of similar inputs by encouraging the encoder to commit to specific codebook entries. The combination of Morphology, Frequency, Demography, Codebook, and Commitment losses enables the model to capture comprehensive features of ECG signals, leading to improved anomaly detection capabilities. These results validate the design of our loss function and underscore the importance of each term in capturing different aspects of the ECG data.
V Discussion
Social Impacts
@@ V Discussion
VI Conclusion
In this paper, we proposed AnyECG, a foundation model family that learns universal embeddings through a two-stage self-supervised pre-training on seven diverse ECG datasets. AnyECG effectively handles the heterogeneity of ECG data through the design of a novel Rhythm Quantizer, which includes a rhythm codebook and a multi-view synergistic decoder to learn representations from different proxy tasks. Additionally, masked modeling in the second-stage pre-training plays a crucial role in enabling effective representation learning of both temporal and lead features of ECG signals. We validated various sizes of AnyECG models on multiple downstream tasks, including anomaly detection, arrhythmia detection, ECG lead generation, and ultra-long ECG signal recognition. Our experiments demonstrate that AnyECG outperforms all state-of-the-art methods in their respective fields, highlighting its effectiveness and versatility in ECG signal analysis.